# Unlocking the Potential of Model Calibration in Federated Learning

### v2

@@ B.3 Mitigating Over/Under-Confidence
In addition to the over-confidence examples discussed in Section 5.2, we also observe cases of under-confidence in FL algorithms. Figure 5 presents the reliability histograms for non-IID ($\alpha=0.5$) FedAvg using the MNIST dataset, where the results show a gap that exceeds the diagonal line, indicating under-confidence issues. While over-confidence is often viewed as more problematic in deep neural networks for real-world applications, under-confidence can also pose challenges in specific scenarios. For example, in a medical diagnosis scenario, a threshold of 0.95 for negative predictions means that any instance classified as negative with a probability $\leq$ 0.95 requires manual review. An underconfident model may classify an instance as negative with a probability of 0.51, suggesting a true probability of correctness that is at least 0.51, yet this does not clarify whether it surpasses the threshold. This leads to compulsory manual inspection regardless of the actual correctness, thus increasing the workload. Among all the figures, we see that our proposed method effectively calibrates FedAvg and exhibits a smaller gap (red region), demonstrating improved performance over other approaches.
@@ B.4 Comparison with FedCal
This section presents additional experimental results and comparisons with FedCal Peng et al. [2024] from Section 5.2. Table 15 provides complete comparisons across different datasets. The results in Table 15 indicate that our proposed NUCFL outperforms FedCal. FedCal improves local client calibration using a local scaler design and aggregates these scalers to create a global scaler for the global model. However, relying solely on an aggregated scaler from local clients can lower global ECE but may neglect crucial interactions between global and local calibration needs. Our method accounts for these interactions, mitigating biases toward local heterogeneity and more effectively meeting global calibration requirements.
Dataset Calibration Method Non-IID FedAvg ($\alpha=0.5$)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
MNIST UnCal. 96.52 1.59 0.60
FedCal 96.52 0.94 0.41
NUCFL(DCA+L-CKA) 98.97 0.48 0.16
FEMNIST UnCal. 91.98 4.38 1.83
FedCal 91.98 3.92 1.56
NUCFL(DCA+L-CKA) 91.89 3.44 1.39
CIFAR-10 UnCal. 75.33 9.69 2.31
FedCal 75.33 7.11 1.90
NUCFL(DCA+L-CKA) 76.96 4.76 1.22
CIFAT-100 UnCal. 72.77 9.80 3.21
FedCal 72.77 5.97 3.03
NUCFL(DCA+L-CKA) 73.73 5.14 2.87
Table 15: Comparison with SOTA baseline on different datasets.

### v3

@@ B.3 Mitigating Over/Under-Confidence
Figure 7: Reliability diagrams for non-IID FedAvg ($\alpha=0.5$) using the FEMNIST dataset.
In addition to the over-confidence examples discussed in Section 5.2, we also observe cases of under-confidence in FL algorithms. Figure 6 and  7 presents the reliability histograms for non-IID ($\alpha=0.5$) FedAvg using the MNIST and FEMNIST datasets, where the results show a gap that exceeds the diagonal line, indicating under-confidence issues. While over-confidence is often viewed as more problematic in deep neural networks for real-world applications, under-confidence can also pose challenges in specific scenarios. For example, in a medical diagnosis scenario, a threshold of 0.95 for negative predictions means that any instance classified as negative with a probability $\leq$ 0.95 requires manual review. An under-confident model may classify an instance as negative with a probability of 0.51, suggesting a true probability of correctness that is at least 0.51, yet this does not clarify whether it surpasses the threshold. This leads to compulsory manual inspection regardless of the actual correctness, thus increasing the workload. Among all the figures, we see that our proposed method effectively calibrates FL and exhibits a smaller gap (red region), demonstrating improved performance over other approaches.
@@ B.4 Comparison with FedCal
This section presents additional experimental results and comparisons with FedCal (Peng et al., 2024) from Section 5.2. Table 23 provides complete comparisons across different datasets. The results in Table 23 indicate that our proposed NUCFL outperforms FedCal. FedCal improves local client calibration using a local scaler design and aggregates these scalers to create a global scaler for the global model. However, relying solely on an aggregated scaler from local clients can lower global ECE but may neglect crucial interactions between global and local calibration needs. Our method accounts for these interactions, mitigating biases toward local heterogeneity and more effectively meeting global calibration requirements.
Dataset Calibration Method IID Non-IID ($\alpha=0.5$) Non-IID ($\alpha=0.1$) Non-IID ($\alpha=0.05$)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
MNIST UnCal. 96.20 0.88 0.29 95.02 1.89 0.65 92.68 2.96 1.04 87.51 6.22 1.83
FedCal 96.20 0.69 0.24 95.02 0.94 0.41 92.68 2.01 0.91 87.51 6.00 1.77
NUCFL(DCA+L-CKA) 96.90 0.44 0.16 95.97 0.78 0.31 92.41 1.32 0.66 87.70 5.20 1.59
FEMNIST UnCal. 92.48 4.10 1.66 90.95 4.17 1.66 88.91 5.58 2.83 86.98 8.78 3.49
FedCal 92.48 3.92 1.66 90.95 4.02 1.62 88.91 5.01 2.62 86.98 8.00 3.42
NUCFL(DCA+L-CKA) 92.39 3.49 1.46 91.63 3.52 1.47 89.89 4.49 1.77 86.89 7.84 3.27
CIFAR-10 UnCal. 80.23 7.60 2.70 78.83 9.69 3.50 64.86 12.62 4.42 59.84 14.30 4.60
FedCal 80.23 7.55 2.70 78.83 8.27 3.19 64.86 11.39 4.40 59.84 14.07 4.58
NUCFL(DCA+L-CKA) 81.05 5.77 1.75 78.66 7.06 2.60 64.77 8.43 3.12 60.30 10.62 3.92
CIFAR-100 UnCal. 65.19 7.61 3.25 61.34 10.52 3.61 58.34 12.02 3.80 56.77 14.80 4.77
FedCal 65.19 7.01 3.20 61.34 8.80 3.49 58.34 12.00 3.80 56.77 13.39 4.56
NUCFL(DCA+L-CKA) 65.43 5.98 2.99 62.05 6.14 3.07 60.73 9.14 3.50 57.03 10.24 3.52
Table 23: Comparison with SOTA baseline (FedCal) on different datasets using FedAvg under different data distribution. The average ACC and calibration error are reported. Our method outperforms the baseline in terms of calibration error and sometimes in average accuracy as well.
B.5 More Difficult Scenario
In addition to evaluating our method in a cross-silo scenario where all clients participate simultaneously, we also assess its performance in a cross-device scenario with partial participation. To accomplish this, we divided the CIFAR-100 training set among $M=100$ clients, each with a distinct data distribution. We then implemented FedAvg using ResNet-34, setting the client participation ratio to 10% for each round. This experiment aims to provide deeper insights into how our method performs under more practical conditions, where participation rates are typically low. Table 24 shows that in this practical scenario, our NUCFL continues to demonstrate its benefits. While the global model may not fully represent the global distribution due to the bias toward clients participating in the previous round, our approach can effectively minimize calibration bias towards current participants.
Calibration Method IID Non-IID ($\alpha=0.5$) Non-IID ($\alpha=0.1$) Non-IID ($\alpha=0.05$)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
UnCal. 57.33 10.21 4.39 54.99 11.49 4.46 51.34 14.02 4.70 49.33 18.80 5.97
DCA 57.61 8.21 3.00 55.10 9.28 3.41 51.24 11.02 3.61 49.29 14.13 4.76
MDCA 57.49 9.40 3.52 55.23 10.33 3.64 51.63 12.39 3.82 49.61 14.66 4.81
FedCal 57.33 9.33 3.49 54.99 9.49 3.46 51.34 10.93 3.60 49.33 13.91 4.74
NUCFL(DCA+L-CKA) 57.55 8.00 2.96 55.23 9.08 3.37 51.93 10.67 3.57 49.29 12.30 3.86
NUCFL(MDCA+L-CKA) 57.21 8.47 3.17 55.39 9.44 3.46 51.49 10.02 3.51 49.69 12.79 4.06
Table 24: Comparison under a partial participants scenario (10% clients in each round) using FedAvg on the CIFAR-100 dataset, where representative calibration methods are selected for comparison. Improved calibration error shows that our method significantly enhances model calibration.
Calibration Method FedSpeed (Sun et al., 2023) FedSAM (Qu et al., 2022) FedMR  (Hu et al., 2023) FedCross  (Hu et al., 2022)
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
UnCal. 65.10 15.33 7.11 64.23 13.95 6.92 64.95 13.18 6.70 65.30 15.81 7.19
FedCal 65.10 12.95 6.53 64.23 12.08 6.48 64.95 12.85 6.49 65.30 14.39 6.98
NUCFL(DCA+L-CKA) 65.17 11.15 5.92 64.52 10.22 5.55 64.95 12.11 6.47 65.33 13.01 6.53
Table 25: Comparison using other FL optimization methods shows that our method can adapt to any FL algorithm and improve calibration error.
Calibration Method local epoch $E=1$ local epoch $E=5$ (default) local epoch $E=10$
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
UnCal. 59.23 8.92 3.47 61.34 10.52 3.61 62.00 11.39 3.98
NUCFL(DCA+L-CKA) 59.47 6.40 3.29 62.05 6.14 3.07 62.05 6.08 3.00
Table 26: Calibration performance with different local epochs using FedAvg under non-IID conditions ($\alpha=0.5$ on CIFAR-100 dataset.
B.6 Trade-off between Accuracy and Reliability
During this study, we observed two types of trade-offs between accuracy and reliability: the Calibration Method Trade-Off and the FL Algorithm Trade-Off.
Calibration Method Trade-Off. In a centralized setup, there are two types of calibration methods, train-time calibration and post-hoc calibration. Post-hoc calibration, which is often used for centralized models, adjusts the calibration error without impacting accuracy. However, as shown in Table 6, when we examine train-time calibration methods in a centralized context, we observe that in over 25% of cases, these methods improve calibration at the cost of reduced accuracy, unlike the stable post-hoc methods. But given that train-time methods are more practical for FL due to the lack of post-hoc dataset, we take this concern into consideration. We found that existing train-time calibration methods tend to degrade accuracy in FL settings, although they vary in their effectiveness at reducing calibration error. This observation further guided our goal to design a calibration method that could improve reliability without compromising accuracy.
FL Algorithm Trade-Off. Another trade-off we identified is between accuracy and reliability across different FL algorithms. While advanced FL algorithms, such as FedDyn and FedNova, can achieve higher accuracy than naive FedAvg, they also tend to introduce greater calibration error. For example, in Table 1, the accuracy (ACC) for FedAvg is 61.34, while FedDyn and FedNova reach 62.39 and 63.01, respectively. However, their calibration error also increases from 10.52 in FedAvg to 12.53 in FedDyn and 11.45 in FedNova. With the help of our method applied to FedDyn and FedNova, we improve their accuracies to 62.84 and 63.17, respectively, while significantly reducing their calibration error to 10.24 and 8.01—both lower than the original FedAvg calibration error of 10.52. This trend is also found in other datasets and data distributions, as shown in our result tables in Appendix B.2. This shows our method’s effectiveness in improving reliability without sacrificing accuracy across various FL algorithms.
B.7 Comparison with FL Optimization Methods Using “Calibration”
While searching for related work, we found that some FL studies Zhang et al. (2022); Luo et al. (2021) use the term “calibration.” Although these works do not focus on “model calibration” as discussed in this paper, we consider them advanced FL optimization methods and evaluate their performance in terms of model calibration. Specifically, we evaluate both the uncalibrated versions of FedLC Zhang et al. (2022) (with $\tau=1$) and CCVR Luo et al. (2021) (on FedAvg) as well as their calibrated counterparts by integrating our proposed NUCFL. We follow the same experimental setup as described in Table 1, utilizing the CIFAR-100 dataset under a non-IID ($\alpha=0.5$) data distribution scenario. The results are provided in Table 27.
Comparing uncalibrated CCVR and FedLC with FedAvg, we observe an increase in accuracy alongside a rise in calibration error. This indicates that these algorithms are advanced FL optimization methods designed primarily to improve accuracy rather than address model calibration. Since we view CCVR and FedLC as advanced FL optimization methods, we can apply NUCFL to them. Comparing the original versions with their NUCFL-calibrated counterparts, we observe improvements in both accuracy and calibration error. This demonstrates the adaptability of our method and its compatibility with any FL algorithm.
Calibration Method FedAvg CCVR FedLC
Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$ Acc $\uparrow$ ECE $\downarrow$ SCE $\downarrow$
UnCal. 61.34 10.52 3.61 63.01 13.69 4.99 62.73 14.58 6.11
NUCFL(DCA+L-CKA) 62.05 6.14 3.07 63.05 9.28 3.53 62.77 10.98 3.79
Table 27: Calibration performance of FL algorithms incorporating “calibration.”
Algorithm 2 Applying NUCFL to FL
2:for each round $t=1,2,...,T$ do
3:   Server sends ${w}^{(t-1)}$ to all clients.
4:   for each client $m\in M$ do
5:      Initialize local model ${w}^{(t,0)}_{m}\leftarrow{w}^{(t-1)}$
6:      for each epoch $e=1,2,...,E$ do
7:         Get the gradient for each client: $\delta_{m}^{(t,e)}={w}^{(t-1)}-{w}^{(t,e)}_{m}$
8:         Calculate the similarity with the previous accumulated gradient: $\beta_{m}=\texttt{sim}(\delta^{(t-1)},\delta_{m}^{(t,e)}),$
9:         Decide calibration loss function $\ell_{cal}$ based on DCA ( 6) or MDCA ( 7).
10:         Each client obtains calibrated local loss $\mathcal{L}_{m}^{cal}({w}_{m}^{(t,e)})$ based on (5).
11:         ${w}^{(t,e)}_{m}\leftarrow\texttt{ClientOPT}({w}^{(t,e-1)}_{m},\mathcal{L}_{m}^{cal}({w}_{m}^{(t,e)}))$
12:      end for
13:      ${w}^{(t,E)}_{m}$ denotes the result after performing $E$ epochs of local updates.
14:      Client sends $\delta^{(t)}_{m}={w}^{(t-1)}-{w}^{(t,E)}_{m}$ to the server after local training.
15:   end for
16:   Server computes aggregate update $\delta^{(t)}=\sum_{m\in M}\frac{|\mathcal{D}_{m}|}{|\mathcal{D}|}\delta^{(t)}_{m}$
17:   Server updates global model ${w}^{(t)}\leftarrow\texttt{ServerOPT}({w}^{(t-1)},\delta^{(t)})$
18:end for